OpenStudy (anonymous):

what is the derivative of y=2e^3x

OpenStudy (cwrw238):

use product rule

OpenStudy (anonymous):

is it y'=6e^3x

OpenStudy (anonymous):

y = 2e^(3x) y* = 2e^(3x) * 3 y* = 6e^(3x)

OpenStudy (anonymous):

You can't use the product rule because you don't have two factors.

OpenStudy (ash2326):

\[y=2e^{3x}\] \[\frac{d}{dx} y=\frac{d}{dx} 2e^{3x}\] \[\frac{dy}{dx}=2 e^{3x} \frac{d}{dx} 3x\] \[\frac{dy}{dx}=2 e^{3x} \times 3=6e^{3x}\]

OpenStudy (anonymous):

okay i got it, thanks.

OpenStudy (cwrw238):

oh i assumed it was 2e^3 * x

OpenStudy (cwrw238):

rishi - to avoid confusion it might be better if , in an instant like this, you write it as 2 e^(3x)

OpenStudy (anonymous):

ok